Find the product or quotient
38. 1 3/8 divided by 2 5/11
Any help with this equation would be Greatly appreciated

Answer by checkley71(8403)

1 3/8=11/8
2 5/11=27/11
11/8*27/11 CANCELL OUT THE 11s.
27/8 ANSWER.